5.10 Find the following values. Compounding/discounting occurs annually. Do not round intermediate calculations. Round your answers...

5.10

Find the following values. Compounding/discounting occurs annually. Do not round intermediate calculations. Round your answers to the nearest cent.

a. An initial $600 compounded for 10 years at 5%.

$  

b. An initial $600 compounded for 10 years at 10%.

$  

c. The present value of $600 due in 10 years at 5%.

$  

d. The present value of $2,470 due in 10 years at 10% and 5%.

Present value at 10%: $  

Present value at 5%: $  

e. Define present value.

  1. The present value is the value today of a sum of money to be received in the future and in general is less than the future value.
  2. The present value is the value today of a sum of money to be received in the future and in general is greater than the future value.
  3. The present value is the value today of a sum of money to be received in the future and in general is equal to the future value.
  4. The present value is the value in the future of a sum of money to be received today and in general is less than the future value.
  5. The present value is the value in the future of a sum of money to be received today and in general is greater than the future value.

-Select-IIIIIIIVVItem 6

How are present values affected by interest rates?

-Select-Assuming positive interest rates, the present value will increase as the interest rate increases.Assuming positive interest rates, the present value will decrease as the interest rate increases.Assuming positive interest rates, the present value will decrease as the interest rate decreases.Assuming positive interest rates, the present value will not change as the interest rate increases.Assuming positive interest rates, the present value will not change as the interest rate decreases.
Answer #1

a) Future value Ś 977.34 =600*(1+5%)^10 b) Future value $1,556.25 -600*(1+10%)^10 c) Present value Ś 368.35 =600/(1+5%)^10 d)

*Please rate thumbs up